OpenStudy (anonymous):

Find all solutions of the equation 2sin^2 x -cos x =1 in the interval [0,2pi).

OpenStudy (anonymous):

just want to know how to start

jimthompson5910 (jim_thompson5910):

hint: sin^2 = 1 - cos^2

OpenStudy (anonymous):

i still dont get it can you write the first part for me

jimthompson5910 (jim_thompson5910):

2sin^2 x -cos x =1 2(1 - cos^2x) -cos x =1 2 - 2cos^2x -cos x =1 -2cos^2x - cosx + 2 = 1 -2cos^2x - cosx + 2 - 1 = 0 -2cos^2x - cosx + 1 = 0 -2z^2 - z + 1 = 0 now solve for z

OpenStudy (anonymous):

thnks

jimthompson5910 (jim_thompson5910):

you're welcome